OpenStudy (anonymous):

Which pair of fractions are equivalent to 3/5 and 2/7? A. 21/35 and 10/35 B. 10/35 and 7/35 C. 7/35 and 5/35 D. 24/35 and 12/35 @phi @iPurple @iGreen @PaulaLovesSchool13 and @baking4lyf

OpenStudy (igreen):

It wants you to change the denominators to 35. To change 3/5 to a denominator of 35, you have to multiply 7 to the top and the bottom. To change 2/7 to a denominator of 35, you have to multiply 5 to the top and the bottom.

OpenStudy (igreen):

3 * 7 = ? 2 * 5 = ? @MeganNagem2002

OpenStudy (igreen):

@MeganNagem2002 Are you there?

OpenStudy (anonymous):

sorry had to go to store back now though

OpenStudy (igreen):

Okay, can you multiply those?

OpenStudy (anonymous):

yes. 7*5=35 and 2*5=10

OpenStudy (anonymous):

wait its 3*7=21 sorry not paying much attention

OpenStudy (anonymous):

@iGreen are you there?

OpenStudy (igreen):

Yes, so we get: 21/35 and 10/35

OpenStudy (anonymous):

okay thanks medaled